ABCDEF is a regular hexagon vector then what is the vector of AB+AC+AD+AE+AF in terms of AO?

Grade:11

1 Answers

Kuldeep Pal
7 years ago
In Regular Hexagon ABCDEF;
 Let AB = a and BC = b.
since AD is parallel to BC and AD = 2BC. Therefore,
 AD = 2b.
 In Triangle ACD, we have
 AC + CD = AD
 CD = 2b – (a + b) = b - a
CE = CD + DE = b – a + (-a) = b – 2a.
 
 
One thing should keep in mind the BOLD and Capital Letters written are All VECTORS and in Same ORDER as written.